Q. choose the correct option.
Q. Let H(x)=∫_a^x▒f(t)dt, for a≤x≤b,then which of the following is true?
(i) H may not be continuous on [a,b]
(ii) f is continuous at c∈[a,b],then H is differentiable at c.
(iii) H is continuous on [a,b].
a.(i) only
b.(i) and (ii) only
c.(ii) and (iii) only.
d.(iii) only
1
Expert's answer
2019-07-11T09:30:40-0400

The correct answer is b) (i) and (ii) only.

I) If f is not integrable on [a, b] (for example, if f is a Dirichlet function), then H will not be continuous on [a, b]. A Dirichlet function is not Riemann integrable (the upper limit sum=1 and the lower limit sum=0), hence in this case "\int_{a}^{x}f(t)dt" doesn't exist.


II) If f is continuous at c∈[a, b], then

"\frac{H(c+h)-H(c)}{h}=\frac{1}{h}\left(\int\limits_a^{c+h}f(t)dt-\int\limits_a^{c}f(t)dt\right)=\frac{1}{h}\int\limits_{c}^{c+h}f(t)dt"

Assuming that


"\int\limits_{c}^{c+h}f(c)dt=hf(c)"

We will have

"\left|\frac{H(c+h)-H(c)}{h}-f(c)\right|=\left|\frac{1}{h}\int\limits_{c}^{c+h}f(t)dt-f(c)\right|=\left|\frac{1}{h}\int\limits_{c}^{c+h}(f(t)-f(c))dt\right|"

Since f is continuous at c

"\lim_{t\to c}\left(f(t)-f(c)\right)=0"

Therefore

"\forall\varepsilon>0 \exist \delta>0, \forall t: |t-c|\le\delta \implies |f(t)-f(c)|\le \varepsilon"

Hence if

"|h|<\delta"

"\left|\frac{1}{h}\int\limits_{c}^{c+h}(f(t)-f(c))dt\right|\le\left|\frac{1}{h}\right|\left|\int\limits_{c}^{c+h}\varepsilon dt\right|=\left|\frac{1}{h}\right|\varepsilon\left|h\right|=\varepsilon"


which means that if "h \to 0"


"\left|\frac{1}{h}\int\limits_{c}^{c+h}(f(t)-f(c))dt\right|\to 0"

Thus, H is differentiable at c